# Stablecoin Runs and Disclosure Policy

### v1

@@ Abstract
Stablecoins have historically depegged due from par to large sales, possibly of speculative nature, or poor reserve asset quality. Using a global game which addresses both concerns, we show that the selling pressure on stablecoin holders increases in the presence of a large sale. While precise public knowledge reduces (increases) the probability of a run when fundamentals are strong (weak), interestingly, more precise private signals increase (reduce) the probability of a run when fundamentals are strong (weak), potentially explaining the stability of opaque stablecoins. The total run probability can be decomposed into components representing risks from large sales and poor collateral. By analyzing how these risk components vary with respect to information uncertainty and fundamentals, we can split the fundamental space into regions based on the type of risk a stablecoin issuer is more prone to. We suggest testable implications and connect our model’s implications to real-world applications, including depegging events and the no-questions-asked property of money.
@@ 1 Introduction
Motivated by these observations, this paper studies the relationship between information uncertainty and the risk from large sales (possibly caused by speculative attacks) or poor collateral (possibly caused by credit events). As of April 2024, stablecoins have a market capitalization of around \$150 billion. [^7] https://www.coingecko.com/en/categories/stablecoins As they form the backbone of many DeFi services (such as automated market makers or lending protocols) and potentially have spillover effects on traditional financial services, it is crucial to understand this relationship. How information uncertainty affects run risk could guide policymaking and regulations on the disclosure and monitoring of reserves.
To capture both types of risk, we extend a streamlined model of Ahmed, Aldasoro, and Duley (2024) to include an large agent that sells at random. To motivate the seller’s randomness, we note that large sales can occur for a variety of reasons, including speculation and the liquidity needs of a large agent, yet small agents will perceive a degree of induced selling pressure regardless. This can also be viewed as repurposing the model of Corsetti et al. (2002) for stablecoins and relaxing it by making the large agent non-strategic. This relaxation allows us to establish uniqueness results and to derive novel comparative statics results with respect to the uncertainty in information of a stablecoin’s reserve asset quality.
@@
Isolating the different types of risk facing a stablecoin also allows us to partition the fundamental space into regions where the conditional run probabilities have differing behaviors with respect to both types of information uncertainty. This classification could be of possible use in assessing what types of risk a stablecoin is prone to given their quality of fundamentals and level of transparency about reserves. The implications generated by our model can be used to explain the events described earlier, and to provide more general insights about the stability mechanisms and regulatory challenges of stablecoins. We finish by discussing the model’s answer to the question of what qualities make a stablecoin to be perceived as money, using the no-questions-asked framework of Holmström (2015).
@@
After the advent of the first cryptocurrencies in the mid-2010s, stablecoins were created to meet the demand for an on-chain asset with relative price stability. Since then, stablecoins have grown in market capitalization, peaking in early 2022 and hovering around the \$150 billion mark. The uses of stablecoins are numerous, including cross-border funds transfers, stores of values for lending protocols, collateral for trading cryptocurrencies, and a medium of exchange in automated market makers.
@@
Our paper contributes to the literature on stablecoin runs. On the game-theoretic side are the works of Gorton et al. (2022), Ma, Zeng, and Zhang (2023), Bertsch (2023), and Ahmed, Aldasoro, and Duley (2024), which all use global games. Gorton et al. (2022) shows that despite the risk of runs, stablecoins maintain their pegs from token holders lending out coins to leveraged traders. Ma, Zeng, and Zhang (2023) focus on the relationship between stablecoin runs and arbitrage centralization, finding a tradeoff between price stability and financial stability as the concentration of the arbitrageur sector varies. Bertsch (2023) explores how the perceived quality of reserves affects the adoption of stablecoins in the presence of outside options. Ahmed, Aldasoro, and Duley (2024) study the effect that public information has on run probability, and is closely related to our work. Ours differs from their in a few key aspects: they do not include large sellers in their model, and we analyze uncertainty in both public and private information.
Compared to the aforementioned works, this paper breaks new ground by introducing a large seller in the context of stablecoins. Our contributions include analyzing the effect of a large seller on selling pressure, considering the dispersion of private signals as a factor influencing stability, and classifying fundamentals with regards to the individual decomposed risk components. For uncertainty in public information, we provide a slightly different result than Ahmed, Aldasoro, and Duley (2024) due to considering the ex-ante run probability (whose distribution also depends on public information uncertainty).
On the empirical side, see Liu, Makarov, and Schoar (2023) which studies stablecoin flows during the Terra–Luna crash, Lyons and Viswanath-Natraj (March 2023) which finds a correlation between peg efficiency and arbitrage volume, and Liao and Caramichael (2022) which examines the effect of stablecoin activity on the real economy. Catalini et al. offers a survey of the economics of stablecoins.
This paper is also related to the literature on global games pioneered by by Carlsson and van Damme (1993), and first applied to currency attacks by Morris and Shin (1998) and Corsetti et al. (2002). Despite similarity in the reduced-form games between our model and currency attack models, the microfoundations differ. Agents in currency attack games all seek to profit from a large sale, whereas we consider the unique struture of the stablecoin market described above and assume that investors find value in holding stablecoins given issuer solvency. Morris and Shin (1998) explore the behavior of speculators in the presence of a government that can choose to defend the peg. Similar to this paper is the work of Corsetti et al. (2002), studying the behavior of a large speculator and small speculators in simultaneous- and sequential-move global games. As previously mentioned, our model contrasts itself by relaxing the speculator’s behavior to be non-strategic in the sequential-move setting. This allows us to derive comparative statics and risk decomposition results, whereas they provide numerical results and characterization of equilibria in asymptotic regimes of the information structure (for the sequential-move setting).
@@ 2.1 Complete Information Game
As a warm-up, we first consider a game where the reserve asset quality is known by all agents.
@@
Time is indexed by $t\in\{0,1,2,3\}$. In Period 0, Nature realizes the value of $\theta$, which is made common knowledge to all investors. In Period 1, with probability $\pi\in[0,1]$, the large seller sells his entire mass of stablecoins (for exogenous reasons). In Period 2, investors observe whether the large seller’s sells or not, then each decide whether to sell their stablecoin or to continue holding on to it. Holding generates a return of $\eta>0$ conditional on the issuer’s solvency. [^8] Returns to holding could arise from investors participating in activities such as lending or staking. Gorton et al. (2022) and Ma et al. (2023) make similar assumptions. In Period 3, the arbitrageur redeems all stablecoins bought from agents in Periods 1–2, possibly causing a stablecoin run. If a run occurs, then the stablecoin drops in price to 0.
@@
Action $=$ Sell $1$ $1$
Action $=$ Hold $1+\eta$ 00
@@
Observe that holding strictly dominates selling if the issuer is solvent and selling strictly dominates holding if the issuer is insolvent. We have the following cases:
  • Suppose that there is no large sale. If $\theta>1-\delta$, then the issuer is solvent even if all investors sell, so the only equilibrium strategy is for all investors to hold. If $\theta\leq 1-\delta$, then there exists a critical proportion $\lambda^{\star}_{0}=\theta/(1-\delta)$. For $\lambda<\lambda^{\star}_{0}$, any selling investor finds it more profitable to hold; for $\lambda>\lambda^{\star}_{0}$, any holding investor finds it more profitable to sell. It follows that the only equilibrium strategies are (i) for all investors to hold or (ii) for all investors to sell.
@@ 2.2 Global Game
To resolve the issue of multiple equilibria, it is common to modify the game into a global game where each agent in the continuum receives a private signal about $\theta$. The agents and market remain the same. The timeline and utilities are modified as follows.
@@
$\mu_{p}(x_{i})=\frac{\sigma^{2}x_{i}+\sigma_{x}^{2}\mu}{\sigma^{2}+\sigma_{x}^{2}},$
$\sigma_{p}=\frac{\sigma^{2}\sigma_{x}^{2}}{\sigma^{2}+\sigma_{x}^{2}}.$
@@
Conditional on fundamentals $\theta$ and the total mass $\lambda$ of redeeming agents, an investor’s utility remains the same. However, investors now evaluate the expected utility with respect to their posterior over $\theta$ given $\bar{x}_{i}$. Moreover, $\lambda$ may also depend on $\theta$. We work with an investor’s expected payoff differential between selling and holding. For an investor receiving a signal of $x_{i}$, this is
@@
where $F_{\mu,\sigma}$ denotes the distribution of a normal random variable with mean $\mu$ and variance $\sigma^{2}$.
@@
[…]
@@ 3 Results
In the previous section, we determined basic results about equilibrium in the global game and the effect of the large seller on the selling behavior of the investors. We now focus our attention to the probability of a stablecoin run implied by an equilibrium switching strategy. The ex-ante run probability evaluted before Nature draws a value of $\theta$ from the $N(\mu,\sigma^{2})$ distribution, which we will denote by $R$, is thus
@@
This decomposition allows us to conceptualize how much two different types of risk present in the stablecoin ecosystem contribute to the overall run probability of an issuer. Although we will have always have $F_{\mu,\sigma}(\bar{\theta}_{0}(x^{\star}_{0}))<F_{\mu,\sigma}(\bar{\theta}_{1}(x^{\star}_{1}))$, the value of $p$ affects the resulting magnitudes of $R_{0}$ and $R_{1}$, as seen in Figure 1. $R_{0}$ and $R_{1}$ may also have differing comparative statics. Proposition 3.1 demonstrates one such case.
@@
Figure 1: A schematic diagram of collateral risk (red) and large sale risk (blue). The horizontal axis represents the value of fundamentals and the vertical axis represents probabilities.
Proposition 3.1.
@@
We now consider the effect of information uncertainty on the risk components and overall run probability. There are two parameters controlling the uncertainty of information in our model: $\sigma$, the scale of the common prior, and $\sigma_{x}$, the scale of dispersion in private signals. Shocks that are known to or perceived by all investors affect $\sigma$, while factors related to the heterogeneity of investors’ beliefs affect $\sigma_{x}$.
@@
$\mu>\bar{\theta}_{a}(x^{\star}_{a})-\frac{F^{-1}(\eta/(1+\eta))}{2\tau_{x}\sqrt{\tau+\tau_{x}}}\implies\frac{dR_{a}}{d\tau}<0,$
@@
When varying the dispersion of the common prior, Proposition 3.2 reveals that for each risk component, there exists upper and lower thresholds at which the risk component increases in $\tau$ for values of $\theta$ below the lower threshold and decreases in $\tau$ for values of $\theta$ above the upper threshold. In other words, when quality of fundamentals is poor, more perceived uncertainty about the value of fundamentals from the investor’s point of view helps reduce the risk of a run. This is because low uncertainty would concentrate the signals of investors around a poor fundamental quality, meaning that an individual investor assesses that other investors have similar pessimistic beliefs about the state of reserves as themselves. Conversely, when the quality of fundamentals is good, more uncertainty increases the risk of a run. The economic reasoning for this is similar: low uncertainty concentrates the signals of investors around a good fundamental quality, so an investor gauges that other investors have similar optimistic beliefs.
Our result is similar to Proposition 2 of Ahmed, Aldasoro, and Duley (2024). They find that a single threshold delineates where increased uncertainty heightens or reduces run risk, while we find two thresholds with a region of ambiguity in the middle. This is because, based on their proof, the notion of run probability that Ahmed, Aldasoro, and Duley (2024) fixes the precision used in evaluating the run probability given a change of $d\tau$, while we consider the effect that that $d\tau$ has on the prior distribution itself as well. The consequence in our case is an extra term when applying the Chain Rule in computing $dR_{a}/d\tau$ that creates a small interval where the sign of $dR_{a}/\partial\tau$ is ambiguous without further assumptions.
@@
[…]
Figure 2: Schematic diagram of how collateral risk and large sale risk change with respect to the precision of the common prior. The horizontal axis represents the mean $\mu$ of the prior distribution.
@@
$\mu>\bar{\theta}_{a}(x^{\star}_{a})-\frac{(2\tau+\tau_{x})\cdot F^{-1}(\eta/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}}\implies\frac{dR_{a}}{d\tau_{x}}>0,$
$\mu<\bar{\theta}_{a}(x^{\star}_{a})-\frac{(2\tau+\tau_{x})\cdot F^{-1}(\eta/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}}\implies\frac{dR_{a}}{d\tau_{x}}<0.$
We now look at the risk components as the dispersion of private signals given $\theta$ varies. Contrary to Proposition 3.2, Proposition 3.3 states that for each component, there exists a single threshold at which the risk component decreases in $\tau_{x}$ for values of $\theta$ below the threshold and increases in $\tau_{x}$ for values of $\theta$ above the threshold. To elaborate, when the quality of fundamentals is good, more dispersion in private signals reduces the run risk, whereas when fundamentals are poor, more varied private signals increase the risk of run. This seemingly paradoxical implication that more uncertainty improves financial stability in good states can be explained by the connection between global games and risk dominance.
Carlsson and van Damme (1993) show that a risk dominant equilibrium is selected in global games. When the variance of private signals is low, the global game approaches the common knowledge game where there are multiple equilibria, one payoff dominant and one risk dominant. As the variance of private signals increases, the investors become more uncertain about each other’s signals and thus actions, heightening the degree of risk dominance. Intuitively, the “safe” option is to hold when fundamentals are good, and sell when fundamentals are poor. Thus, Propositions 3.2 and 3.3 contrast as the common prior’s variance drives investors’ uncertainty about the stablecoin reserves whereas the private signal’s variance drives investors’ uncertainty about each other’s actions.
[…]
Figure 4: Schematic diagram of how collateral risk and large sale risk change with respect to the precision of private signals. The horizontal axis represents the mean $\mu$ of the prior distribution.
@@
$S_{0}=\frac{R_{0}}{R_{0}+R_{1}},$
$S_{1}=\frac{R_{1}}{R_{0}+R_{1}}.$
As an immediate corollary of Proposition 3.1, we have that the share of large sale risk increases in the size of the large seller. For the regions implied by Propositions 3.2 and 3.3 where one type of risk increases and the other decreases as information is more precise, we immediately deduce that the share of large sale risk decreases in the variance of the common prior for $\mu\in(\bar{\theta}_{0}(x^{\star}_{0})-\frac{F^{-1}(\eta/(1+\eta))}{2\tau_{x}\sqrt{\tau+\tau_{x}}},\bar{\theta}_{1}(x^{\star}_{1})$ and increases in the variance of private signals for $\mu\in(\bar{\theta}_{0}(x^{\star}_{0})-\frac{(2\tau+\tau_{x})\cdot F^{-1}(\eta/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}},\bar{\theta}_{1}(x^{\star}_{1})-\frac{(2\tau+\tau_{x})\cdot F^{-1}(\eta/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}})$. As the average quality of fundamentals itself, captured by $\mu$, changes, we have the following result.
Proposition 3.4.

### v2

# Information Structures in Stablecoin Markets

@@ Abstract
Stablecoins have historically depegged due from par to large sales, possibly of speculative nature, or poor reserve asset quality. Using a global game which addresses both concerns, we show that the selling pressure on stablecoin holders increases in the presence of a large redemption. While precise public knowledge reduces (increases) the probability of a run when fundamentals are strong (weak), interestingly, more precise private signals increase (reduce) the probability of a run when fundamentals are strong (weak), potentially explaining the stability of opaque stablecoins. The total run probability can be decomposed into components representing risks from large sales and poor collateral. By analyzing how these risk components vary with respect to information uncertainty and fundamentals, we can split the fundamental space into regions based on the type of risk a stablecoin issuer is more prone to. We suggest testable implications and connect our model’s implications to real-world applications, including depegging events, stablecoin regulations, and the no-questions-asked property of money.
@@ 1 Introduction
Motivated by these observations, this paper studies the relationship between information uncertainty and the risk from large sales (possibly caused by speculative attacks) or poor collateral (possibly caused by credit events). As of November 2025, stablecoins have a market capitalization of around \$300 billion. [^6] https://www.coingecko.com/en/categories/stablecoins As they form the backbone of many DeFi services (such as automated market makers or lending protocols) and potentially have spillover effects on traditional financial services, it is crucial to understand this relationship. How information uncertainty affects run risk could guide policymaking and regulations on the disclosure and monitoring of reserves. Indeed the implications of our model can be used to analyze the effect of recent stablecoin legislation, such as the GENIUS Act (U.S.) and MiCA regulations (E.U.), that have increased requirements for stablecoin issuers.
To capture both types of risk, we extend a streamlined model of [1] to include an large agent that sells at random. To motivate the seller’s randomness, we note that large sales can occur for a variety of reasons, including speculation and the liquidity needs of a large agent, yet small agents will perceive a degree of induced selling pressure regardless. This can also be viewed as repurposing the model of [4] for stablecoins and relaxing it by making the large agent non-strategic. This relaxation allows us to establish uniqueness results and to derive novel comparative statics results with respect to the uncertainty in information of a stablecoin’s reserve asset quality.
@@
Isolating the different types of risk facing a stablecoin also allows us to partition the fundamental space into regions where the conditional run probabilities have differing behaviors with respect to both types of information uncertainty. This classification could be of possible use in assessing what types of risk a stablecoin is prone to given their quality of fundamentals and level of transparency about reserves. The implications generated by our model can be used to explain the events described earlier, and to provide more general insights about the stability mechanisms and regulatory challenges facing stablecoins.
@@
After the advent of the first cryptocurrencies in the mid-2010s, stablecoins were created to meet the demand for an on-chain asset with relative price stability. Since then, stablecoins have grown in market capitalization, peaking in early 2022 and hovering around \$300 billion as of January 2026. The uses of stablecoins are numerous, including cross-border funds transfers, stores of values for lending protocols, collateral for trading cryptocurrencies, and a medium of exchange in automated market makers.
@@
In July 2025, the United States government passed the GENIUS Act, establishing a comprehensive framework for payment stablecoins. Notably, to be a permitted stablecoin issuer, an entity must maintain full reserve backing of the token with money market assets (e.g. dollars, Treasures, “similarly liquid assets”), publicly disclose their token redemption policy, and publish monthly audits or reports of their reserve composition. The legislation does not provide deposit insurance or Fed backstops for stablecoins, instead placing responsibility on the issuer to manage reserves and disclose information. [^7] See https://www.congress.gov/bill/119th-congress/senate-bill/394 for the full text.
@@
We contribute to the literature on stablecoin runs. On the game-theoretic side are the works of [5], [9], [2], and [1], which all use global games. [5] shows that despite the risk of runs, stablecoins maintain their pegs from token holders lending out coins to leveraged traders. [9] focuses on the relationship between stablecoin runs and arbitrage centralization, finding a tradeoff between price stability and financial stability as the concentration of the arbitrageur sector varies. [2] explores how the perceived quality of reserves affects the adoption of stablecoins in the presence of outside options. [1] study the effect that public information has on run probability, and is closely related to our work. We differ from them in a few key aspects by accounting for large sellers and analyzing uncertainty in both public and private information.
Our contribution is threefold. First, we study the effect of a large seller on selling pressure. Second, we consider the dispersion of private signals as a factor influencing stability. Third, we introduce a risk framework based on our model and use it to evaluate stablecoin fundamentals.
On the empirical side, see [7] which studies stablecoin flows during the Terra–Luna crash, [8] which finds a correlation between peg efficiency and arbitrage volume, and [6] which examines the effect of stablecoin activity on the real economy. Catalini et al. offers a survey of the economics of stablecoins.
This paper is also related to the literature on global games pioneered by by [3], and first applied to currency attacks by [10] and [4]. Despite similarity in the reduced-form games between our model and currency attack models, the microfoundations differ. Agents in currency attack games all seek to profit from a large sale, whereas we consider the unique struture of the stablecoin market described above and assume that investors find value in holding stablecoins given issuer solvency. [10] explore the behavior of speculators in the presence of a government that can choose to defend the peg. Similar to this paper is the work of [4], studying the behavior of a large speculator and small speculators in simultaneous- and sequential-move global games. Our model contrasts itself by relaxing the speculator’s behavior to be non-strategic in the sequential-move setting. This allows us to derive comparative statics and risk decomposition results, whereas they provide numerical results and characterization of equilibria in asymptotic regimes of the information structure (for the sequential-move setting).
@@ 2.1 Complete Information Game
Time is indexed by $t\in\{0,1,2,3\}$. In Period 0, Nature realizes the value of $\theta$, which is made common knowledge to all investors. In Period 1, with probability $\pi\in[0,1]$, the large seller sells his entire mass of stablecoins (for exogenous reasons). In Period 2, investors observe whether the large seller’s sells or not, then each decide whether to sell their stablecoin or to continue holding on to it. Holding generates a return of $\eta>0$ conditional on the issuer’s solvency. [^8] Returns to holding could arise from investors participating in activities such as lending or staking. [5] and [9] make similar assumptions. In Period 3, the arbitrageur redeems all stablecoins bought from agents in Periods 1–2, possibly causing a stablecoin run. If a run occurs, then the stablecoin drops in price to 0.
@@
Action == Sell 11 11
Action == Hold $1+\eta$ 0
@@
Observe that holding strictly dominates selling if the issuer is solvent and selling strictly dominates holding if the issuer is insolvent. Then:
  • Suppose that there is no large sale. If $\theta>1-\delta$, then the issuer is solvent even if all investors sell, so the only equilibrium strategy is for all investors to hold. If $\theta\leq 1-\delta$, then there exists a critical proportion $\lambda^{\star}_{0}=\theta/(1-\delta)$. For $\lambda<\lambda^{\star}_{0}$, selling investors finds it more profitable to hold; for $\lambda>\lambda^{\star}_{0}$, holding investors finds it more profitable to sell. It follows that the only equilibrium strategies are (i) for all investors to hold or (ii) for all investors to sell.
@@ 2.2 Global Game
To resolve the issue of multiple equilibria, we modify the game into a global game where each agent in the continuum receives a private signal about $\theta$. The agents and market remain the same. The timeline and utilities are modified as follows.
@@
$\mu_{p}(x_{i})=\frac{\sigma^{2}x_{i}+\sigma_{x}^{2}\mu}{\sigma^{2}+\sigma_{x}^{2}},\quad\sigma_{p}=\frac{\sigma^{2}\sigma_{x}^{2}}{\sigma^{2}+\sigma_{x}^{2}}.$
@@
Remark 1
A natural modeling question is whether the large seller’s decision should be treated as endogenous rather than exogenously specified as it is in our case. Compared to [4], which considers an endogenous large seller, this model produces the same key insight (of increased selling pressure) without introducing multiplicity of equilibria in the global game. Additionally, it focuses the analysis on the event of a large redemption occurring without committing to a specific motive, which may vary (such as speculation, liquidity needs, risk management). The resulting framework allows for uniqueness, tractability, and interpretable comparative statics.
@@
For fundamentals $\theta$ and the total mass $\lambda$ of redeeming agents, an investor’s utility remains the same. However, they now evaluate the expected utility with respect to their posterior over $\theta$ given $\bar{x}_{i}$. Moreover, $\lambda$ may also depend on $\theta$. We work with an investor’s expected payoff differential between selling and holding. For an investor receiving a signal of $x_{i}$, this is
@@
where $F_{\mu,\sigma}$ is the distribution of a $N(0,\sigma^{2})$ random variable.
@@
$\Delta_{0}(x_{i};\bar{x})=\int_{-\infty}^{\bar{\theta}_{0}(\bar{x})}dF_{\mu_{p}(x_{i}),\sigma_{p}}(\theta)-\int_{\bar{\theta}_{0}(\bar{x})}^{\infty}\eta\,dF_{\mu_{p}(x_{i}),\sigma_{p}}(\theta)$
$\Delta_{1}(x_{i};\bar{x})=\int_{-\infty}^{\bar{\theta}_{1}(\bar{x})}dF_{\mu_{p}(x_{i}),\sigma_{p}}(\theta)-\int_{\bar{\theta}_{1}(\bar{x})}^{\infty}\eta\,dF_{\mu_{p}(x_{i}),\sigma_{p}}(\theta)$
@@ 3 Results
In the previous section, we determined basic results about equilibrium in the global game and the effect of the large seller on the selling behavior of the investors. We now focus our attention to the probability of a stablecoin run implied by an equilibrium switching strategy. The ex-ante run probability $R$ evaluated before Nature draws a value of $\theta$ from the $N(\mu,\sigma^{2})$ distribution is
@@
This decomposition allows us to conceptualize how much two different types of risk present in the stablecoin ecosystem contribute to the overall run probability of an issuer. Although we will have always have $F_{\mu,\sigma}(\bar{\theta}_{0}(x^{\star}_{0}))<F_{\mu,\sigma}(\bar{\theta}_{1}(x^{\star}_{1}))$, the value of $p$ affects the magnitudes of $R_{0}$ and $R_{1}$, as seen in Figure 1.
@@
Figure 1: A diagram of collateral risk (red) and large sale risk (blue). The horizontal axis represents the value of fundamentals and the vertical axis represents probabilities.
Proposition 3.1
@@
We now consider the effect of information uncertainty on the risk components and overall run probability. Two parameters control the uncertainty of information in our model: the scale of the common prior $\sigma$ and the scale of dispersion in private signals $\sigma_{x}$. Shocks that are known to all investors affect $\sigma$, while factors related to the heterogeneity of investors’ beliefs affect $\sigma_{x}$.
@@
$\mu>\bar{\theta}_{a}(x^{\star}_{a})+\frac{F^{-1}(1/(1+\eta))}{2\tau_{x}\sqrt{\tau+\tau_{x}}}\implies\frac{dR_{a}}{d\tau}<0,$
@@
When varying the dispersion of the common prior, Proposition 3.2 reveals that for each risk component, there exists upper and lower thresholds at which the risk component increases in $\tau$ for values of $\theta$ below the lower threshold and decreases in $\tau$ for values of $\theta$ above the upper threshold. In other words, when quality of fundamentals is poor, more perceived uncertainty about the value of fundamentals from the investor’s point of view helps reduce the risk of a run. This is because low uncertainty concentrates investors’ signals around a poor fundamental quality, meaning that an individual investor assesses that other investors have similar pessimistic beliefs about the state of reserves as themselves. Conversely, when the quality of fundamentals is good, more uncertainty increases the risk of a run. The economic reasoning for this is similar: low uncertainty concentrates the signals of investors around a good fundamental quality, so an investor gauges that other investors have similar optimistic beliefs.
$\frac{\partial R_{1}}{\partial\tau}$$\frac{\partial R_{0}}{\partial\tau}$$\overline{\theta}_{0}\!\left(x_{0}^{\star}\right)$-$\overline{\theta}_{1}\!\left(x_{1}^{\star}\right)$$\overline{\theta}_{0}\!\left(x_{0}^{\star}\right)+\frac{F^{-1}(1/(1+\eta))}{2\tau_{x}\sqrt{\tau+\tau_{x}}}$$\overline{\theta}_{1}\!\left(x_{1}^{\star}\right)+\frac{F^{-1}(1/(1+\eta))}{2\tau_{x}\sqrt{\tau+\tau_{x}}}$
Figure 2: Diagram of how collateral risk and large sale risk change with respect to the precision of the common prior. The horizontal axis represents the prior mean $\mu$.
@@
$\mu>\bar{\theta}_{a}(x^{\star}_{a})+\frac{(2\tau+\tau_{x})\cdot F^{-1}(1/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}}\implies\frac{dR_{a}}{d\tau_{x}}>0,$
$\mu<\bar{\theta}_{a}(x^{\star}_{a})+\frac{(2\tau+\tau_{x})\cdot F^{-1}(1/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}}\implies\frac{dR_{a}}{d\tau_{x}}<0.$
We now look at the risk components as private signal dispersion varies. Contrary to Proposition 3.2, Proposition 3.3 states that for each component, there exists a single threshold at which the risk component decreases in $\tau_{x}$ for values of $\theta$ below the threshold and increases in $\tau_{x}$ for values of $\theta$ above the threshold. To elaborate, when the quality of fundamentals is good, more dispersion in private signals reduces the run risk, whereas when fundamentals are poor, more varied private signals increase the risk of run. This seemingly paradoxical implication that more uncertainty improves financial stability in good states can be explained by the connection between global games and risk dominance.
[3] show that a risk dominant equilibrium is selected in global games. When the variance of private signals is low, the global game approaches the common knowledge game where there are multiple equilibria, one payoff dominant and one risk dominant. As the variance of private signals increases, the investors become more uncertain about each other’s signals and thus actions, heightening the degree of risk dominance. Intuitively, the “safe” option is to hold when fundamentals are good, and sell when fundamentals are poor. Thus, Propositions 3.2 and 3.3 contrast as the common prior’s variance drives investors’ uncertainty about the stablecoin reserves whereas the private signal’s variance drives investors’ uncertainty about each other’s actions.
$\frac{\partial R_{1}}{\partial\tau_{x}}$$\frac{\partial R_{0}}{\partial\tau_{x}}$-$\overline{\theta}_{0}\left(x_{0}^{\star}\right)+\frac{(2\tau+\tau_{x})\cdot F^{-1}(1/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}}$$\overline{\theta}_{1}\left(x_{1}^{\star}\right)+\frac{(2\tau+\tau_{x})\cdot F^{-1}(1/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}}$++
Figure 4: Diagram of how collateral risk and large sale risk change with respect to the precision of private signals. The horizontal axis represents the prior mean $\mu$.
@@
$S_{0}=\frac{R_{0}}{R_{0}+R_{1}},\quad S_{1}=\frac{R_{1}}{R_{0}+R_{1}}.$
As an immediate corollary of Proposition 3.1, we have that the share of large sale risk increases in the size of the large seller. For the regions implied by Propositions 3.2 and 3.3 where one type of risk increases and the other decreases as information is more precise, we can deduce that the share of large sale risk decreases in the variance of the common prior and increases in the variance of private signals for an intermediate range of $\mu$. As the average quality of fundamentals itself, captured by $\mu$, changes, we have the following result.
Proposition 3.4